solve by the addition method 3x-2y=-1 2x-3y=-4
Answer by checkley77(12569)

3x-2y=-1 multiply by -2
2x-3y=-4 multiply by 3 & add
-6x+4y=2
6x^2-9y=-12
-----------------
-5y=-10
y=-10/-5
y=2 ans.
3x-2*2=-1
3x-4=-1
3x=-1+4
3x=3
x=1 ans.
Proof:
2*1-3*2=-4
2-6=-4
-4=-4